OpenStudy (dariyana):

Given the triangle and equation below, what is the value of x? sec∠C - cot∠A = 3x

OpenStudy (dariyana):

7-\sqrt{13}\div18\]

OpenStudy (dariyana):

6-\sqrt{13}\div7\]

OpenStudy (dariyana):

3\sqrt{13}\div13\]

OpenStudy (dariyana):

sqrt{13}\div3\]

OpenStudy (btaylor):

sec C = 7/6 cot A = sqrt13/6 sec C - cot A = (7-sqrt{13})/6 divide that by 3, and you get \[x = \frac{7-\sqrt{13}}{18}\]

OpenStudy (dariyana):

THANK YOU SO MUCH YOU ARE MY SAVER @BTaylor

OpenStudy (btaylor):

you're welcome.